OpenStudy (anonymous):

(5k/m^2)^3 AND 9y^8/15xy^3 Help me! SHow work, too

OpenStudy (anonymous):

cloverpetal see ur previous post..i gave the answers

OpenStudy (anonymous):

(5k/m^2)^3 = (5^3*k^3)/(m^(2*3) =125k^3/m^6

OpenStudy (anonymous):

THANKS YOU SO MUCH.... what about the second one?

OpenStudy (anonymous):

(9y^8)/(15xy^3)=(3y^8)/(5xy^3) =(3y^5)/(5x)

OpenStudy (anonymous):

Oh my goodness, thank you so much!

OpenStudy (anonymous):

THANks so much!!